- Press Release
- Sep 28, 2022
NOAA GOES SATOPS Morning Report 31 October 2000
SATOPS MORNING REPORT – October 31, 2000
No change in the status of the POES spacecraft.
NOAA-16 operations nominal.
A NOAA-15 solar array adjustment was executed without incident at 19:07Z
on Monday October 30, to further shade the IMP/AVHRR. Synch delta
information will be collected while in site of the ground stations, both
before (at 1725Z) and approximately 48 hours after the array movement.
On Wednesday, November 1 at 18:22Z, the AVHRR will be switched from high
to low power mode and further synch delta samples will be collected.
Additional configuration changes to investigate the AVHRR performance are
The NOAA-15 HIRS Filter Wheel Housing Heater was turned on yesterday
at 22:24Z due to continuing degradation of filter wheel performance.
This activity was coordinated with NASA and represents the final active
measure to mitigate poor filter wheel performance. Long term performance
is not expected to improve, but engineers continue to monitor HIRS performance
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
At 2032Z on Monday, October 30, a GOES-8 frame break was observed and
attributed to loss of signal sync.
GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode
at ~105 deg W.
GOES operations were nominal over the past 24 hours.
Questions or comments to firstname.lastname@example.org